Diffusion Morning Report 12/06/06

Ted Berg tberg at stanford.edu
Wed Dec 6 06:34:39 PST 2006


Good Morning All,
    Looking good this morning. Epi leak checking is moving along 
slowly-FS has been requested. Adding a 2GB hard drive to ebeam so it 
hopefully will not
fill up so fast. Transfering PLC power to a more dedicated circuit. AMAT 
process in yesterday to wrap up tier 2. Moving on to tier 3. TEL 
received first quote for demoing existing piping. No Other issues at 
this time.  LTO going down next Wed. for pull and clean. Several LP logs 
not updated or not updated correctly.

Diffusion 	  	  	  	  	 
  	Date: 	12/6/06 	THICKNESS-Change at 	  	 
Tube Name/Number 	Status 	Comments 	  	  	 
Tylan 1 	Up 	  	  	  	 
Tylan 2 	Up 	  	  	  	 
Tylan 3 	Up 	  	  	  	 
Tylan 4 	Up 	  	  	  	 
Tylan 5-BBr3 	Up 	  	  	  	 
Tylan 6-POCl3 	Up 	  	  	  	 
Tylan 7-TEOS2 	Up 	  	9.25/30 	  	 
Tylan 8-SiGe 	Up 	  	9.13/30 	  	 
Tylan 9-Poly 	Up 	  	18.78/30 	  	 
Tylan 10-Nitride 	Up 	  	27.49/30 	  	 
Tylan 11-BPSG-LTO 	UP 	  	28.06/30 	  	 
Tylan 12-FGA 	Up 	  	  	  	 
  	  	  	  	  	 
Thermco 1 	Up 	  	  	  	 
Thermco 2 	Up 	  	  	  	 
Thermco Poly 	Up 	  	  	  	 
Thermco Nitride 	Up 	  	  	  	 
Epi 	down 	leak testing 	  	  	 
  	
	
	
	
	
Burn Box 	up 	  	  	  	 
ZYGO 	Up 	  	  	  	 
  	up 	  	   	  	 
GAAS 22 	  	  	  	  	 
GAAS 23 	UP 	  	  	  	 
SiH4 for furnaces 	up 	  	  	  	 

	  	A on line 540 psi 	 changeover working well  	  	 




-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20061206/fb79b612/attachment.html>


More information about the diffusion mailing list